# BA-ORABE: Blockchain-Based Auditable Registered Attribute-Based Encryption With Reliable Outsourced Decryption

### v1

@@ Abstract
Attribute-based encryption (ABE) is a generalization of public-key encryption that enables fine-grained access control in cloud services. Recently, Hohenberger et al. (Eurocrypt 2023) introduced the notion of registered ABE, which is an ABE scheme without a trusted central authority. Instead, users generate their own public/secret keys and then register their keys and attributes with a key curator. The key curator is a transparent and untrusted entity and its behavior needs to be audited for malicious registration. In addition, pairing-based registered ABE still suffers the heavy decryption overhead like ABE. A general approach to address this issue is to outsource decryption to a decryption cloud service (DCS). In this work, we propose BA-ORABE, the first fully auditable registered ABE with reliable outsource decryption scheme based on blockchain. First, we utilize a verifiable tag mechanism to achieve verifiability of ciphertext transformation, and the exemptibility which enables the honest DCS to escape from wrong claims is guaranteed by zero knowledge fraud proof under optimistic assumption. Additionally, our system achieves fairness and decentralized outsourcing to protect the interests of all parties and the registration and outsourcing process are transparent and fully auditable through blockchain. Finally, we give formal security analysis and implement and evaluate our scheme on Ethereum to demonstrate its feasibility and efficiency.
@@ 1 Introduction
The swift progression of cloud computing technologies has led to a significant transformation in data sharing landscape. Nowadays, both individuals and enterprises are increasingly reliant on cloud storage platforms for uploading and disseminating data. However, given that cloud service providers (CSPs) may not always be deemed fully reliable and considering the high sensitivity of much of the data stored in the cloud, ensuring access control within cloud storage is paramount.
@@
However, Attribute-Based Encryption(ABE) upends the traditional trust model of public-key encryption by relying on a single trusted authority that maintains a long-term master secret key to issue decryption keys[2]. If the central authority is compromised and its secret key is obtained by an adversary, the adversary would be able to decrypt all ciphertexts within the system. The proposal of decentralized ABE still relies on multiple authorities, and therefore, it fails to completely address this issue[3]. To address the key-escrow problem, Hohenberger et al.[4] proposed a registered ABE scheme without any authority that allows users to generate secret keys on their own and then register the associated public key with a key curator along with their attributes. The key curator combines the public keys of different users into a single master public key. Users may need to obtain auxiliary decryption keys from the key curator, which they then combine with their own secret keys to decrypt. The key curator is transparent and does not maintain secrets, allowing for public auditing, as both registration and update algorithms are deterministic. Although a malicious key curator could register attributes that allow the decryption of all ciphertexts, such behavior can always be detected by an external auditor. By utilizing a decentralized, immutable ledger, blockchain[5] can provide auditable records of all operations performed by the key curator. This ensures that any malicious behavior, such as the registration of unauthorized attributes or wrong computation, can be detected by public verifier. The tamper-resistant nature of blockchain enhances the integrity of the system by enabling the verification and traceability of all interactions without reliance on a central authority.
@@ 1.1 Motivation and Contribution
While registered ABE tackles the key-escrow problem upon traditional ABE by removing central authority, it requires robust auditing to ensure the integrity of key registration and update processes, as the transparent key curator still plays a critical role in the system. Additionally, it still suffers from high decryption overhead like ABE, making it unsuitable for lightweight devices. Many verifiable OABE schemes have been proposed to address this issue on ABE and several OABE schemes[16, 17] achieve exemptibility but outsource heavy computation to blockchain, leading to high gas fees and performance issues. To the best of our knowledge, there is no outsource decryption scheme currently existing for registered ABE. This gap motivates our work to develop the first outsourced decryption scheme for registered ABE, which aims to reduce decryption overhead while ensuring security, verifiability, exemptibility, and fairness based on blockchain. In summary, the contributions of this paper are threefold.
@@
  • We implement a concrete auditable registered ABE scheme with reliable and payable outsourced decryption and evaluate feasibility and performance on Ethereum.
@@ 1.2 Related Work
1.2.1 Outsource Decryption Attribute Based Encryption
Attribute-Based Encryption (ABE) was first introduced as fuzzy identity-based encryption by Sahai and Waters in [18]. ABE schemes are generally classified into two categories: Ciphertext-Policy ABE (CP-ABE) and Key-Policy ABE (KP-ABE) [6], depending on whether the access policy is embedded in the ciphertext or the user’s private key. A user can decrypt only if the set of attributes satisfies the access policy of ciphertext. However, the decryption in the ABE is too expensive, and this problem is especially acute for resource limited devices such as mobile devices, which greatly hinders its real-world adoption. To mitigate this issue, Green et al. [9] proposed outsourcing the majority of the decryption to third party without revealing the plaintext. The pioneer work[9] can append a hash of ciphertext randomness as tag to check the correctness of transformation. Consequently, Lai et al. [11] extend the original ABE model with outsourced decryption to achieve verifiability. Their approach involved attaching an additional ciphertext of a random message and a tag to each original ciphertext which double the size of the original ABE ciphertext and the transformation cost. Based on the scheme [11], Qin et al. [13] and Lin et al. [14] aimed to enhance the efficiency of OABE. Qin et al.[13] managed to reduce the ciphertext size on both the server and client sides by half and achieved faster decryption than the scheme proposed in [11]. They also introduced a method to transform any ABE scheme with outsourced decryption into one that supports verifiable outsourced decryption. Lin et al. [14] proposed a more efficient and generic construction of ABE with verifiable outsourced decryption based on an attribute-based key encapsulation mechanism, a symmetric-key encryption scheme, and a commitment scheme. Furthermore, Li et al. [12] introduced the notion of ABE scheme that support both outsourced encryption and decryption, but only supports the threshold access policy. Following their work, Ma et al. [19] extended it to accommodate any monotonic access policy and established a more stronger security framework compared to [11]. None of the aforementioned schemes can achieve both exemptibility and fairness. To tackle this issue, Cui et al. [16] introduced an payable OABE scheme that leverages blockchain technology to achieve exemptibility and fairness. However, it still introduces redundant information. Ge et al. [17] eliminated redundant information and mitigated heavy computation problem of smart contracts through decomposing to pieces of meta computations, but their solution still results in high gas consumption and performance limitations. Furthermore, due to the lack of support for pairing computation in native Ethereum[20], this scheme is incompatible with Ethereum platform.
1.2.2 Registered Attribute Based Encryption
@@ 1.3 Organization
The rest of this paper is organized as follows. In Section 2, we briefly revisit the definitions that are relevant to this paper. In Section 3, we describe the system overview of our scheme. In Section 4, we present generic construction of slotted ORABE and ORABE scheme using blockchain and zero-knowledge fraud proof. In section 5, we give formal security analysis of our scheme. In Section 6, we implement the proposed concrete scheme on Ethereum to evaluate its performance. Finally, this paper is concluded in Section 7.
@@ 3.1 System Architecture
  • DO encrypts sensitive data using RABE under Key Encapsulation Mechanism(KEM) and sends ciphertext to IPFS.
@@
Note that the dispute phase and public verifier are needed only when one player crashes or attempts to cheat under our optimistic assumption.
@@ 3.2 Threat Model
3.3 Security Definition
  • Static Security. Let $\prod_{sORABE}=(\textbf{Setup},\textbf{KeyGen},\textbf{IsValid},\textbf{Aggregate},\textbf{Encrypt},$ Transform, $\textbf{Decrypt}_{out})$ be a slotted outsource decryption registered ABE scheme with attribute universe $\mathcal{U}$ and policy space $\mathcal{P}$. For a security parameter $\lambda$, an adversary $\mathcal{A}$, and a bit $b\in\{0,1\}$, we define the following game $Game_{sec}$ between a challenger C and an adversary $\mathcal{A}$:
@@ 3.3 Security Definition
    A Registered ABE scheme with outsourced decryption is secure if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
@@
  • Verifiability of outsourced decryption. Verifiability guarantees that the correctness of transform ciphertext can be efficiently checked. It is described by a game $Game_{ver}$ between a challenger C and an adversary $\mathcal{A}$. $Game_{ver}$ consists of the following steps:
    • Setup: C runs Setup gives crs to $\mathcal{A}$.
    • Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to C, C runs RegPK and gives mpk to $\mathcal{A}$.
    • Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},sk^{*},CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}\}$, where S* satisfies $(M^{*},\rho^{*})$, $CT^{*}$ is an original ciphertext, $sk^{*}$ is the secret key of $\mathcal{A}$ and $CT_{1}^{*^{\prime}},CT_{2}^{*^{\prime}}$ are two transformed ciphertexts of $CT^{*}$. $\mathcal{A}$ wins the game if:
      $Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq\perp\cap$
      $Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})\neq\perp\cap$
      $Decrypt_{out}(sk^{*},CT_{1}^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk^{*},CT_{2}^{*^{\prime}},CT^{*})$
    The advantage of the adversary in $Game_{ver}$ is defined as $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{ver}]$.
    Definition 3.2.
    A Registered ABE scheme with outsourced decryption is verifiable if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
    $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{ver}]\leq negl(\lambda)$
  • Exemptibility of outsourced decryption. Exemptibility guarantees that the user can not falsely accuse DCS of returning wrong transform ciphertext if it is not in the case. It is described by a game $Game_{exe}$ between a challenger C and an adversary $\mathcal{A}$. $Game_{exe}$ consists of the following steps:
    • Setup: C runs Setup and gives crs to $\mathcal{A}$.
    • Challenge $\mathcal{A}$ runs keygen generate $(sk^{*},pk^{*})$ and register (S*,pk*) to C, C runs RegPK and gives mpk and aux to $\mathcal{A}$.
    • Outputs $\mathcal{A}$ outputs an access structure $(M^{*},\rho^{*})$ and a tuple $\{CT^{*},CT^{*^{\prime}},hsk*,(sk_{1}^{*},\pi_{1}),(sk_{2}^{*},\pi_{2}),T_{i}^{*}\}$, where S* satisfies $(M^{*},\rho^{*})$, $CT^{*}$ is an original ciphertext, $CT^{*^{\prime}}$ is the transform ciphertext of $CT^{*}$, $hsk*,sk_{1}^{*},sk_{2}^{*}$ are the helper decryption key and secret keys, $T_{i}^{*}=g^{sk^{*}}$ is part of public key of $\mathcal{A}$ and $\pi_{1},\pi_{2}$ are the corresponding NIZK proof for secret keys. $\mathcal{A}$ wins the game if:
      $hsk^{*}\leftarrow Update(crs,aux,pk^{*})\cap$
      $CT^{*^{\prime}}\leftarrow Transform(hsk^{*},CT^{*})\cap$
      $1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{1}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{1}^{*}},g,T_{i}^{*}))\cap$
      $1\leftarrow\prod_{DLEQ-NIZK}.\textbf{Verify}(\pi_{2}^{\prime},(C_{2}^{\prime},C_{2}^{\prime sk_{2}^{*}},g,T_{i}^{*}))\cap$
      $Decrypt_{out}(sk_{1}^{*},CT^{*^{\prime}},CT^{*})\neq Decrypt_{out}(sk_{2}^{*},CT^{*^{\prime}},CT^{*})$
    The advantage of the adversary in $Game_{exe}$ is defined as $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{exe}]$.
    Definition 3.3.
    A Registered ABE scheme with outsourced decryption is exempt if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game,
    $\textbf{Pr}[\mathcal{A}\quad\text{wins}\quad Game_{exe}]\leq negl(\lambda)$
@@ 4 Construction
We first present slotted registered ABE and registered ABE scheme with outsourced decryption based on registered ABE scheme in [8]. Then we describe a NIZK fraud proof mechanism under optimistic assumption to achieve exemptibility. Finally, we present the concrete smart contract design for our payable outsourced decryption scheme.
@@ 4.1 Verifiable Outsourced Decryption Slotted Registered ABE
    Finally, it selects three hash functions $H_{0},H_{1},H_{2}$ and outputs the master public key mpk and the slot-specific helper decryption keys $hsk_{i}$ where
@@
  • $\textbf{Transform}(hsk,ct)$rightarrow ct’: On input the ciphertext $ct$ and the helper decryption key $hsk$. The transform algorithm proceeds as follows:
    • If the set of attributes $S_{i}$ is not authorized by $(\textbf{M},\rho)$, then the transform algorithm outputs $\perp$.
    • Otherwise, let $I=\{k\in[K]:\rho(k)\in S_{i}\}$ be the indices of the rows of M associated with the attributes $S_{i}\subseteq\mathcal{U}_{\lambda}$.
    • Let $\textbf{M}_{S_{i}}$ be the matrix formed by taking the subset of rows in M indexed by $I$. Since $S_{i}$ is authorized, let $\omega_{S_{i}}\in\mathbb{Z}_{p}^{|I|}$ be a vector such that $\omega_{S_{i}}^{\top}\textbf{M}_{S_{i}}=\textbf{e}_{1}^{\top}$.
    • Then compute:
@@ 4.3 NIZK Fraud Proof Mechanism
Unfortunately, this verify algorithm can not be executed by Ethereum smart contract because it does not support $\mathbb{G}_{t}$ computation of any elliptic curve now. So, we introduce a semi-honest and transparent public verifier to verify the fraud proof and publish verification result on blockchain. Ge et al. [17] added precompile contract to native Ethereum to support pairing and $\mathbb{G}_{t}$ computation on their own blockchain. If reduce the reliance on native Ethereum security, we can integrate their approach to eliminate the role for semi-honest public verifier and enable the verification of fraud proof directly through smart contracts. In this work, we make a trade-off and prioritize the security of native Ethereum. Note that in the registered ABE scheme, this role can also be played by KC, who is semi-honest and transparent.

### v2

@@ Abstract
Attribute-based encryption (ABE) is a generalization of public-key encryption that enables fine-grained access control in cloud services. Recently, Hohenberger et al. (Eurocrypt 2023) introduced the notion of registered ABE, which is an ABE scheme without a trusted central authority. Instead, users generate their own public/secret keys and then register their keys and attributes with a key curator. The key curator is a transparent and untrusted entity and its behavior needs to be audited for malicious registration. In addition, pairing-based registered ABE still suffers the heavy decryption overhead like ABE. A general approach to address this issue is to outsource decryption to a decryption cloud service (DCS). In this work, we propose BA-ORABE, the first fully auditable registered ABE with reliable outsourced decryption scheme based on blockchain. First, we utilize a verifiable tag mechanism to achieve verifiability of ciphertext transformation, and the exemptibility which enables the honest DCS to escape from wrong claims is guaranteed by zero knowledge fraud proof under optimistic assumption. Additionally, our system achieves fairness and decentralized outsourcing to protect the interests of all parties and the registration and outsourcing process are transparent and fully auditable through blockchain. Finally, we give security analysis, implement and evaluate our scheme on Ethereum to demonstrate its feasibility and efficiency, and show its advantages in real application of decentralized finance.
@@ 1 Introduction
The swift progression of cloud computing technologies has led to a significant transformation in data sharing landscape. Nowadays, both individuals and enterprises are increasingly reliant on cloud storage platforms for uploading and disseminating data. However, since cloud service providers(CSPs) may not always be perceived as fully reliable, and given the high sensitivity of much of the data stored in the cloud, ensuring access control in cloud storage is crucial.
@@
However, Attribute-Based Encryption(ABE) upends the traditional trust model of public-key encryption by relying on a single trusted authority that maintains a long-term master secret key to issue decryption keys[2]. If the central authority is compromised and its secret key is obtained by an adversary, the adversary would be able to decrypt all ciphertexts within the system. The proposal of decentralized ABE still relies on multiple authorities, and therefore, it fails to completely address this issue[3]. To address the key-escrow problem, Hohenberger et al.[4] proposed a registered ABE scheme without any authority that allows users to generate secret keys on their own and then register the associated public key with a key curator along with their attributes. The key curator combines the public keys of different users into a single master public key. Users may need to obtain auxiliary decryption keys from the key curator, which they then combine with their own secret keys to decrypt. The key curator is transparent and does not maintain secrets, allowing for public auditing, as both registration and update algorithms are deterministic. Although a malicious key curator could register attributes that allow the decryption of all ciphertexts, such behavior can always be detected by an external auditor. By utilizing a decentralized and immutable ledger, blockchain[5] can provide auditable records of all operations performed by the key curator. This ensures that any malicious behavior, such as the registration of unauthorized attributes or wrong computation, can be detected by public verifier. The tamper-resistant nature of blockchain enhances the integrity of the system by enabling the verification and traceability of all interactions without reliance on a central authority.
@@ 1.1 Motivation and Contribution
While registered ABE tackles the key-escrow problem upon traditional ABE by removing central authority, it requires robust auditing to ensure the integrity of key registration and update processes, as the transparent key curator still plays a critical role in the system. Additionally, it still suffers from high decryption overhead like ABE, making it unsuitable for lightweight devices. Many verifiable OABE schemes have been proposed to address this issue on ABE and several OABE schemes[16, 17] achieve exemptibility but outsource heavy computation to blockchain, leading to high gas fees and performance issues. To the best of our knowledge, there is no outsourced decryption scheme currently existing for registered ABE. This gap motivates our work to develop the first outsourced decryption scheme for registered ABE, which aims to reduce decryption overhead while ensuring security, verifiability, exemptibility, and fairness based on blockchain. In summary, the contributions of this paper are threefold.
@@
  • We implement a concrete auditable registered ABE scheme with reliable and payable outsourced decryption, evaluate its feasibility and performance on Ethereum, and show its application in decentralized finance(DeFi).
@@ 1.2 Related Work
1.2.1 Outsourced Decryption of Attribute-Based Encryption
Attribute-Based Encryption (ABE) was first introduced as fuzzy identity-based encryption by Sahai and Waters in [18]. ABE schemes are generally classified into two categories: Ciphertext-Policy ABE (CP-ABE) and Key-Policy ABE (KP-ABE) [6], depending on whether the access policy is embedded in the ciphertext or the user’s private key. A user can decrypt only if the set of attributes satisfies the access policy of ciphertext. However, the decryption in the ABE is too expensive, and this problem is especially acute for resource limited devices such as mobile devices, which greatly hinders its real-world adoption. To mitigate this issue, Green et al. [9] proposed outsourcing the majority of the decryption to third party without revealing the plaintext. The pioneer work[9] can append a hash of ciphertext randomness as tag to check the correctness of transformation. Consequently, Lai et al. [11] extend the original ABE model with outsourced decryption to achieve verifiability. Their approach involved attaching an additional ciphertext of a random message and a tag to each original ciphertext which double the size of the original ABE ciphertext and the transformation cost. Based on the scheme [11], Qin et al. [13] and Lin et al. [14] aimed to enhance the efficiency of OABE. Qin et al.[13] managed to reduce the ciphertext size on both the server and client sides by half and achieved faster decryption than the scheme proposed in [11]. They also introduced a method to transform any ABE scheme with outsourced decryption into one that supports verifiable outsourced decryption. Lin et al. [14] proposed a more efficient and generic construction of ABE with verifiable outsourced decryption based on an attribute-based key encapsulation mechanism, a symmetric-key encryption scheme, and a commitment scheme. Furthermore, Li et al. [12] introduced the notion of ABE scheme that support both outsourced encryption and decryption, but only supports the threshold access policy. Following their work, Ma et al. [19] extended it to accommodate any monotonic access policy and established a more stronger security framework compared to [11]. But the exemptibility they claimed is not achieved since the user may reveal an incorrect private key and accuse the decryption cloud server falsely. None of the aforementioned schemes can achieve both exemptibility and fairness. To tackle this issue, Cui et al. [16] introduced an payable OABE scheme that leverages blockchain technology to achieve exemptibility and fairness. However, it still introduces redundant information. Ge et al. [17] eliminated redundant information and mitigated heavy computation problem of smart contracts through decomposing to pieces of meta computations, but their solution still results in high gas consumption and performance limitations. Furthermore, due to the lack of support for pairing computation in native Ethereum[20], this scheme is incompatible with Ethereum platform.
1.2.2 Registered Attribute-Based Encryption
@@
1.2.3 Applications of Optimistic Fraud Proof
The fraud proof operates under an optimistic assumption, defaulting to accept statement unless proven otherwise. It includes a challenge window, during which challengers may initiate disputes to demonstrate the statement’s invalidity. In contrast, the validity proof offers evidence confirming the correctness of the statement, with acceptance granted only when its accuracy is definitively established under a pessimistic assumption. Optimistic rollups[27],[28],[29],[30],[31],[32], a key solution for Ethereum’s scalability, assumes that most transactions are valid. This assumption facilitates faster and more efficient transaction processing. To guard against invalid transactions, fraud proofs are utilized to identify and dispute inconsistencies. Many projects[27],[28] implement multi-round fraud proofs using the interactive bisection protocol[33], while others[31],[32] adopt zero-knowledge proofs to enable single-round fraud proofs, thereby reducing the challenge window. Beyond optimistic rollups, this fraud-proof mechanism is widely applied in other blockchain contexts, such as the Lightning Network[34], Plasma[35], and distributed key generation protocols[36].
@@ 1.3 Organization
The rest of this paper is organized as follows. In Section 2, we briefly revisit the definitions that are relevant to this paper. In Section 3, we describe the system overview of our scheme. In Section 4, we present generic construction of slotted ORABE and ORABE scheme using blockchain and zero-knowledge fraud proof. In section 5, we give formal security analysis of our scheme. In Section 6, we implement the proposed concrete scheme on Ethereum to evaluate its performance. In Section 7, we describe the application of our scheme in DeFi. Finally, this paper is concluded in Section 8.
@@ 3.1 System Architecture
  • DO encrypts sensitive data using registered ABE under Key Encapsulation Mechanism(KEM) and sends ciphertext to IPFS.
@@
Note that the optimistic assumption supposes that both parties are honest unless proven otherwise. The dispute phase and public verifier are needed only when one player crashes or attempts to cheat under our optimistic assumption.
@@ 3.2 Threat Model
3.3 Security Requirement
  • Static Security. Let $\prod_{sORABE}=(\textbf{Setup},\textbf{KeyGen},\textbf{IsValid},\textbf{Aggregate},\textbf{Encrypt},$ Transform, $\textbf{Decrypt}_{out})$ be a slotted outsourced decryption registered ABE scheme with attribute universe $\mathcal{U}$ and policy space $\mathcal{P}$. For a security parameter $\lambda$, an adversary $\mathcal{A}$, and a bit $b\in\{0,1\}$, we define the following game $Game_{sec}$ between a challenger C and an adversary $\mathcal{A}$:
@@ 3.3 Security Requirement
    A registered ABE scheme with outsourced decryption is secure if all probabilistic polynomial time adversaries have at most a negligible advantage in the above game:
@@
  • Verifiability of outsourced decryption. The verifiability guarantees that the correctness of transform ciphertext can be efficiently checked.
  • Exemptibility of outsourced decryption. The exemptibility guarantees that the user can not falsely accuse DCS of returning wrong transform ciphertext through fraud proof if it is not in the case.
  • Fairness of outsourced decryption. The fairness property in our work guarantees that:
    • DCS gets paid if and only if it has returned a correct result.
    • DU must pay for the correct result and can challenge wrong result through fraud proof.
    Fairness incentivizes each party to perform the protocol honestly.
@@ 4 Construction
We first present slotted registered ABE and registered ABE scheme with outsourced decryption based on registered ABE scheme in [8]. Then we describe a NIZK fraud proof mechanism under optimistic assumption to achieve exemptibility. Finally, we present the concrete smart contract design for our auditable and fair outsourced decryption scheme.
@@ 4.1 Verifiable Outsourced Decryption Slotted Registered ABE
    Finally, it selects three collision-resistant hash functions $H_{0},H_{1},H_{2}$ and outputs the master public key mpk and the slot-specific helper decryption keys $hsk_{i}$ where
@@
  • $\textbf{Transform}(hsk,ct)$rightarrow ct’: On input the ciphertext $ct$ and the helper decryption key $hsk$. If the set of attributes $S_{i}$ is not authorized by $(\textbf{M},\rho)$, then the transform algorithm outputs $\perp$. Otherwise, let $I=\{k\in[K]:\rho(k)\in S_{i}\}$ be the indices of the rows of M associated with the attributes $S_{i}\subseteq\mathcal{U}_{\lambda}$. Let $\textbf{M}_{S_{i}}$ be the matrix formed by taking the subset of rows in M indexed by $I$. Since $S_{i}$ is authorized, let $\omega_{S_{i}}\in\mathbb{Z}_{p}^{|I|}$ be a vector such that $\omega_{S_{i}}^{\top}\textbf{M}_{S_{i}}=\textbf{e}_{1}^{\top}$. Then compute:
@@ 4.3 NIZK Fraud Proof Mechanism
Unfortunately, this verify algorithm can not be executed by Ethereum smart contract because it does not support $\mathbb{G}_{t}$ computation of any elliptic curve now. So, we introduce a semi-honest and transparent public verifier to verify the fraud proof and publish verification result on blockchain. Ge et al. [17] added precompile contract to native Ethereum to support pairing and $\mathbb{G}_{t}$ computation on their own blockchain. If reduce the reliance on native Ethereum security, we can integrate their approach to eliminate the role for semi-honest public verifier and enable the verification of fraud proof directly through smart contracts. In this work, we make a trade-off and prioritize the security of native Ethereum. Note that in the registered ABE scheme, this public verifier role can also be played by KC, who is semi-honest and transparent.
Figure 3: Contract design for BA-ORABE.